OpenStudy (anonymous):

If the radius of a pizza is 20 cm, what is the central angle in degrees that gives one person 30 cm of crust?

OpenStudy (johan14th):

Do you know the equation to find the angle?

OpenStudy (anonymous):

\[\theta=\frac{ arc length }{ radius }\]

OpenStudy (johan14th):

So that would be

OpenStudy (johan14th):

theta = (30/20)

OpenStudy (anonymous):

yes \[\frac{ 3}{ 2 }\]

OpenStudy (anonymous):

in degrees i got 270